OpenStudy (anonymous):

Given the first term and the common ratio of a geometric sequence, find the first five terms and write the explicit formula. a1 = 1, r = 2

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

That's the formula for the nth term of tier 1 geometric sequence

OpenStudy (anonymous):

whats tn?

OpenStudy (anonymous):

the nth term

OpenStudy (anonymous):

\[T_{n} = a _{1} \times r ^{n - 1}\]

OpenStudy (anonymous):

the first term is 1 the '`common ratio " is 2 so the second term is \(2\times 1=2\) ad the third term is \(2\times 2=4\) and so on

OpenStudy (anonymous):

just keep multiplying by 2 \[1,2,4,8,16,32,...\]

OpenStudy (anonymous):

@satellite73 are those the first five terms

OpenStudy (kropot72):

The first term is a1 = 1 The second term a2 is the first term multiplied by the common ratio = 1 * r = 1 * 2 = 2 The third term a3 is the first term multiplied by (the common ratio) squared: a3 = 1 * r * r = 1 * 2 * 2 = 4 The fouth term a4 is the first term multiplied by (the common ratio) cubed: a4 = 1 * r * r * r = 1 * 2 * 2 * 2 = 8 The fifth term a3 is the first term multiplied by (the common ratio) to the power of 4: a5 = 1 * 2 * 2 * 2 * 2 = you can calculate.

OpenStudy (anonymous):

actually those are the first six terms, yes

OpenStudy (anonymous):

so what's the explicit formula?

OpenStudy (anonymous):

the 6th term is \(2^5=32\) ad i general the "nth" term is \(\large2^{n-1}\)

OpenStudy (anonymous):

@satellite73 so is 2^(n-1) the explicit formula?

OpenStudy (kropot72):

The general formula is \[n ^{th}\ term=a _{1}\times r ^{n-1}\]